1. Insurance The 30th September 2017 Financial Accounts reported a prepayment of 1500 pounds. In the year of 30th September 2018 records show payments made of 1800 pounds. The 1800 pounds includes a prepayment for 2019 of 1200 pounds. what is the expense to be reported in 2018? 2. Legal fee The 30th September 2017 Financial Accounts reported an accrual of 2500 pounds. In the year to 30th September 2018 records show payments made of 20000 pounds. Notes suggest there is a invoice due for services received worth 3500 pounds. What is the expense to be reported in 2018? 3. Cost of sales The 30th September 2017 Financial Accounts reported inventories of 49500 pounds. In the year of 30th September 2018 records show purchases of 198500 pounds. On 30th September 2018 inventory was 38000 pounds. What is the expense to be reported in 2018? 4. Depreciation-plant&equipment 1st October 2017 ABC purhased equipment worth 125000 pounds. Useful economic life is expected to be 5 years,and residual value 20000 pounds. The straight line method to be used. What is the expense to be reported in 2018?
Answer #1

1) Insurance expense in 2018 = 1500+1800-1200 = $2100 Pounds

2) Legal fees expense in 2018 = 20000-2500+3500 = $21000 Pounds

3) Cost of sales = 49500+198500-38000 = $210000 Pounds

4) Depreciation expense = (125000-20000/5) = $21000 Pounds